# A Generative Approach to Control Complex Physical Systems

### v1

@@ E.4 Training, Inference, and Evaluation
Appendix F 1D Baselines
F.1 PID
@@
As PID itself is a SISO control method, ANN-PID uses a neural network to get multiple sets of PID parameters to do multiple SISO PID controls for MIMO control in the Burgers’ equation. But here, ANN-PID requires the dimensions of inputs and outputs to be the same, so it can only cope with full observation, full control control problems, and partial observation, partial control control problems. Besides, the ANN-PID controller has 2 training setups, including directly interacting with the solver and interacting with the 1D surrogate model in Appendix H.
F.2 SAC
@@
F.3 Supervised Learning
The paper [23] proposes a supervised-learning-based control algorithm that takes a neural operator as a surrogate model to solve control problems. It contains two stages. In the first stage, we take a neural operator to learn the physical constraint as Appendix H. The three CNNs respectively reconstruct $u$, reconstruct $\mathbf{w}$ and learn the transition from $u_{t}$ to $u_{t+1}$. More details are in Appendix H. In the second stage, these three neural networks are used as surrogate models to calculate the gradient of the objective function with respect to the control input. We consider the control $\mathbf{w}$ as a learnable parameter and update it with the gradient.
@@ F.5 BPPO
Appendix G 2D Baselines
G.1 MPC and SL
Model Predictive Control (MPC) [59] is a control strategy that solves an optimization problem repeatedly to determine the optimal control inputs for a dynamic system. It operates over a finite prediction horizon, optimizing a cost function and applying only the first control action. In the 2D jellyfish movement control problem, MPC uses the control sequences $\mathbf{w}$ and the fluid states as internal state variables. Without the need to train a control agent model, MPC relies on 2D surrogate models mentioned in Appendix H to estimate future states based on the current state and control. We use backpropagation to compute the gradient, update the control action sequences, and optimize the control objective $\mathcal{J}$ in Eq. (19). For every time step, we get the optimized sequences from this time step forward in this way, and only the first control sequence of the optimized control sequences is applied. Compared with MPC, the Supervised learning (SL) method [23] only optimizes the entire control sequences and employs the entire sequences.
@@ G.1 MPC and SL
G.2 SAC
For the 2D case, the algorithm and basic architecture of SAC are the same as the 1D case in Appendix F.2. When designing the reward function for the 2D jellyfish movement control, we find the periodic condition of the opening angle curves is hard to constrain. So to satisfy the periodic condition better, we include the distance between $\mathbf{w}_{t}$ of every time step $t$ and $\mathbf{w}_{0}$. Also, we both consider the squared and absolute error of $(\mathbf{w}_{t}-\mathbf{w}_{0})$ since they respectively constrain the periodic condition when $(\mathbf{w}_{t}-\mathbf{w}_{0})$ is large and small. As a result, the reward function of time step $t$, force $F_{t}$, opening angle $(\mathbf{w}_{t-1},\mathbf{w}_{t})$ and condition angle $\mathbf{w}_{0}$ is defined as
@@ G.2 SAC
where $\lambda_{1}$, $\lambda_{2}$ are weights of different terms. We take the Adam optimizer [28] to update the weights of networks and the temperature parameter as in the 1D experiment. The hyperparameters are reported in Table 12. In particular, we take the best checkpoint to evaluate the final performance, thus the actual number of training episodes for each setting ranges from about 100 to about 300.
@@
G.3 BC
For the 2D task, the algorithm and basic architecture of BC are the same as the 1D case in Appendix F.4. In Behavior Cloning, the policy network is trained using supervised learning techniques, where the objective is to minimize the difference between the predicted actions and the actions taken by the non-expert in the training data. The loss function is typically the mean squared error between the predicted actions and actions in training data. The training data, consisting of state-action pairs, is collected from non-expert random demonstrations and stored in a dataset. In the inference, we evaluated the same objective function as the SAC. The detailed values of hyperparameters are provided in Table 13.
@@
Hyperparameters for 2D Jellyfish movement control:
@@
G.4 BPPO
For the 2D task, the algorithm and basic architecture of BPPO are the same as the 1D case in Appendix F.5. When designing the reward function for the 2D jellyfish movement control, we find the periodic condition of the opening angle curves is hard to constrain. So to satisfy the periodic condition better, we include the distance between $\mathbf{w}_{t}$ of every time step $t$ and $\mathbf{w}_{0}$. Also, we both consider the squared and absolute error of $(\mathbf{w}_{t}-\mathbf{w}_{0})$ since they respectively constrain the periodic condition when $(\mathbf{w}_{t}-\mathbf{w}_{0})$ is large and small. As a result, the reward function of time step $t$, force $F_{t}$, opening angle $(\mathbf{w}_{t-1},\mathbf{w}_{t})$ and condition angle $\mathbf{w}_{0}$ is defined as
@@ G.4 BPPO
Hyperparameters for 2D jellyfish movement control:
@@
Appendix H Surrogate Models
H.1 1D Surrogate Model
@@ Appendix H Surrogate Models
H.3 2D Boundary Mask and Offsets Updater
@@
H.5 DiffPhyCon for 1D Burgers’ Equation
@@
As can be observed from Figure 12, there is a clear correlation between $\mathcal{J}$ and $\mathcal{E}$. This is because, the higher the prediction error is, the less accurate the solution to the inverse control problem would be. The correlation proves that our DiffPhyCon learns the dynamics of the system (i.e., the dependency of $\mathbf{u}$ on $\mathbf{w}$), based on which the control sequence is generated. In partial observation settings, the randomness is larger than in the full observation setting, which demonstrates the challenge of controlling partially observed systems. The reference line denotes $\mathcal{J}=\mathcal{E}$ where the control objective completely depends on the prediction accuracy. In Figure 12 (b), samples cluster tightly around the line, which implies DiffPhyCon can find near-optimal control sequence based on the predicted trajectory in the corresponding level of prediction accuracy. This ability is remarkable in that DiffPhyCon does not rely on explicit gradient descent like those in the adjoint method and SL.
H.6 2D Jellyfish Movement
In this subsection, we record extensive results of 2D jellyfish movement control. To further compare baselines with DiffPhyCon in more settings, we change $\zeta$ in $\mathcal{J}$ and results of $\zeta=500$, $\zeta=1000$ and $\zeta=2000$ are listed in Table 16 and Table 17. From the results, it is easy to observe that DiffPhyCon outperforms other baselines in diverse settings with different objectives.
We also present more simulation results of our method in Figure 13. Each line represents an example from the test dataset. We plot five snapshots of boundary and fluid field for each example.
@@
Figure 13: More examples of 2D simulation visualizations controlled by our method.
H.7 Myopic failure modes of SAC
@@
Appendix I Efficiency Comparison
We test the inference (control) efficiency of DiffPhyCon and baselines on both 1D Burgers’ equation and 2D jellyfish movement tasks on a single Tesla-V100 GPU with 8 CPUs. On the 1D Burgers’ equation task, we test the efficiency of partial observation and full control; on the 2D jellyfish movement task, we test the efficiency of full observation and full control. The average inference times over all test samples are listed in Table 19 and Table 20, respectively. These results reveal that although DiffPhyCon is not as efficient as reinforcement learning methods like SAC, BC, and BPPO, it has competitive efficiency compared to SL, and MPC. In particular, by using the fast sampling method DDIM [62], DiffPhyCon could be accelerated significantly by reducing the number of samplings.
Table 19: Efficiency comparison on 1D Burgers’ equation on a Tesla-V100 GPU with 8 CPUs.
Methods Time (s)
DiffPhyCon-lite 21.13
DiffPhyCon 58.97
DiffPhyCon-DDIM (8 sampling steps) 0.53
PID 3.61
SL 74.85
SAC 0.11
BC 1.22
BPPO 0.82
Table 20: Efficiency comparison on 2D jellyfish movement on a Tesla-V100 GPU with 8 CPUs.
Methods Time (s)
DiffPhyCon 252.2
DiffPhyCon-DDIM (8 sampling steps) 12.6
MPC 1401.7
SL 133.5
SAC 0.2
BC 0.99
BPPO 1.08
Appendix J Effect of Hyperparameters
J.1 Effect of $\gamma$
@@ Appendix I Efficiency Comparison
We tested different scheduling of the prior reweighting $\gamma$ and found their performance similar. Finally, we use the same cosine schedule as that in the DDPM noise for clarity. Our experiment results also revealed that the performance of DiffPhyCon in 1D Burgers equation control is insensitive to the prior reweighting intensity $\gamma$ as we reported in Table 1. Besides, we conducted a more comprehensive experiment on the prior reweighting intensity $\gamma$ in all three settings (FO-PC; PO-FC; PO-PC) as in Table 22, 23, and 24. The results show that prior reweighting does not significantly impact the performance of DiffPhyCon in 1D Burgers control task, where State MSE denotes the deviation between generated $u$ and the ground truth $u_{\text{gt}}$ given by the numerical simulator based on generated $f$.
$\gamma$ 0.0 0.3 0.5 0.7 1.0
$\mathcal{J}_{\text{actual}}$ 0.00038 0.00037 0.00037 0.00037 0.00037
State MSE 0.00063 0.00043 0.00034 0.00028 0.00025
$\gamma$ 0.0 0.3 0.5 0.7 1.0
$\mathcal{J}_{\text{actual}}$ 0.01159 0.01155 0.01152 0.01148 0.01139
State MSE 0.02643 0.02642 0.02642 0.02641 0.02639
$\gamma$ 0.0 0.3 0.5 0.7 1.0
$\mathcal{J}_{\text{actual}}$ 0.00493 0.00493 0.00493 0.00493 0.00494
State MSE 0.00675 0.00667 0.00665 0.00664 0.00663
@@
In such a setting, varying $\lambda$ changes how the generated samples are shifted towards lower energy regions, but an excessively large $\lambda$ causes the sample to deviate from the physically feasible region. Results in Table 26, 27, and 28 show that the behavior of DiffPhyCon under different $\lambda$ is expected. Note that “physically feasible” samples here are samples with lower $\mathcal{J}_{\text{actual}}$.
@@
2D jellyfish movement control. In our jellyfish movement control task, we use explicit guidance with $\lambda$ involved. Since diffusion models denoise gradually, similar to the hyperparameter $\gamma$, we use a varying sequence of $\lambda=$ $\lambda_{k}$$=\lambda_{0}\times$ $\beta_{k}$, where $\beta_{k}$ is the schedule of variances of noise in DDPM [2] increasing monotonically from $\beta_{1}=0.0003$ to $\beta_{K}=1$. The motivation of this schedule is that the confidence of gradient estimation becomes stronger along with the denoising process from $k=K$ to $k=1$. The results are shown in the following table. We can observe that as $\lambda_{0}$ (which determines the $\lambda$ schedule) increases from 0 to 0.6, the performance of DiffPhyCon improves consistently. Compared to Table 3, DiffPhyCon always outperforms baselines for a wide range of $0.1\leq\lambda_{0}\leq 0.5$. When $\lambda_{0}\geq 0.6$, the results are no longer valid since the generated control sequences may be unfeasible. This reflects that $\lambda$ should not be overly large to make it safe, which is a consistent conclusion with our previous analysis.
@@
Appendix K Extensions to Finer-grained Jellyfish Movement Control Task
@@
K.1 Experimental Setting
Similar to previous setups in Appendix E, this extension aims to control the movement of a flapping jellyfish with two wings in a 2D fluid field where fluid flows with constant initial speed. However, in this task, we parametrize the jellyfish’s boundary as the coordinate change ($\Delta x$, $\Delta y$) for each cell within the boundary, which serves as the control sequence $\mathbf{w}$. Undoubtedly, this is a high-dimensional and complex control task. Firstly, the control sequence $\mathbf{w}$ is elevated to three dimensions. In addition to ensuring optimization towards the control objective, maintaining consistency in the movement of different cells within the jellyfish boundary is also crucial. Bearing this in mind, we have chosen to employ DiffConPDE-lite for this experimental setup to minimize potential disruptions to the jellyfish boundary. Regarding the PDE state, the experiment adopts the full observation setup. All other settings remain the same with Appendix E.
K.2 Data Generation
@@
K.4 Training, Inference, and Evaluation
@@
Appendix L Future Work
Moving forward, we plan to expedite the inference process by drawing insights from relevant prior works, such as [57, 40, 32]. Additionally, we aim to build upon our existing trajectory knowledge to generalize future trajectories. Furthermore, we intend to incorporate solver interactions into our training methodology, enhancing the model’s adaptability and effectiveness.
Appendix M Social Impact Statements

### v3

@@ F.4 Training, Inference, and Evaluation
Appendix G Additional Details for 2D Smoke Control
G.1 Dataset Preparation
Figure 14: 2D smoke indirect control task. There are seven exits in total and the top middle one is the target exit (a). The control signals are only allowed to apply to peripheral regions (b). The control objective is to minimize the proportion of smoke failing to pass through the target exit.
Table 8: Obstacle Positions
Category Position
Bottom (8:56,8:9)
Left (8:9,8:12)
(8:9,20:28)
(8:9,36:56)
Right (56:57,8:12)
(56:57,20:28)
(56:57,36:56)
Up (8:12,56:57)
(20:28,56:57)
(36:44,56:57)
(52:56,56:57)
Inside Obstacles (24:25,32:40)
(24:25,48:56)
(40:41,32:40)
(40:41,48:56)
(20:44, 20:21)
Table 9: Absorb Area
Category Absorb Area
Left (0:8, 11:21)
(0:8, 27:37)
Right (56:64, 11:21)
(56:64, 27:37)
Up (11:21, 56:64)
(27:37, 56:64)
(43:53, 56:64)
We use the Phiflow solver [22] to generate the incompressible fluid with an indirect control dataset. The resolution of the 2D flow field is set to be 64 $\times$ 64. The flow field is set to be boundless in Phiflow. We placed obstacles and absorbing areas in the middle of the fluid domain. In Figure 14, we illustrate the locations of the exits, obstacles, and controllable areas. The specific positions of the obstacles and absorbing areas are detailed in Table 9 & Table 9.
At the beginning of each trajectory, we set the horizontal velocity component $v_{x}$ of the entire flow field to zero and the vertical component $v_{y}$ to an upward velocity of 1.0. Additionally, we randomly initialize a square smoke patch with dimensions of $4\times 4$ in the area below the horizontal obstacles, specifically within the coordinates (11:50, 11:14). In the trajectories we designed, the smoke is required to make four turns. We randomly determine the positions where the smoke will turn, and calculate the smoke’s horizontal ($v_{x}$) and vertical ($v_{y}$) velocity components, assuming a constant magnitude of velocity. By adjusting the parameters, we observed that when the $v_{x}$ added at the turning points is sampled from a Gaussian distribution $\mathcal{N}(mv_{x},m^{2}v_{x}^{2}/16)$ and $v_{y}$ is sampled from a Gaussian distribution $\mathcal{N}(6v_{y},9v_{y}^{2}/4)$, where is $m$ sampled from a uniform distribution $U(2,7)$, the success rate of the smoke navigating through is approximately 0.5387% in the training set and 0.5286% in the testing set, which meets our requirements. During the motion of the smoke, at turning points, the velocity in the central region is set to the velocity from the previous moment, while the velocity in the surrounding regions is assigned by randomly sampling from the distributions $\mathcal{N}(v_{turn},{v_{turn}}^{2}/100)$. At non-turning points, the velocity in the central region remains the velocity from the previous moment, but the velocity in the surrounding areas is adjusted to the previous velocity plus noise sampled from $\mathcal{N}(0,0.01)$. This procedure ensures that each control parameter varies at every moment and remains distinct from others at any given time. Such variability is beneficial for the subsequent training and generation of control parameters.
We duplicated the density field into two versions: the original density field and the set-zero density field. The original density field is used for model training and remains unaffected by the absorbing areas. In contrast, the set-zero density field is employed to calculate the amount of smoke passing through each bucket. When set-zero density is present in the absorb areas, we sum up and record this density. Once the recording is complete, we reset the set-zero density in the absorb areas to zero to prevent the double-counting of emitted smoke. Ultimately, we document the quantity of smoke emitted from each bucket at every moment.
Our experimental data records the velocity and density of the entire flow field at each moment, as well as the velocity of the peripheral flow field (control field) after noise addition. We generated 40,000 training trajectories and 500 test trajectories, with each trajectory approximately 10 MB in size. Given a total of 40,500 trajectories, the entire dataset is estimated to be around 400 GB in size.
G.2 Experimental Setting
In smoke control, we only consider the full observation and full control setting. Namely, all fluid features, including smoke density, horizontal and vertical velocities are observable; and all green areas shown in Figure 14 are controllable.
G.3 Model
G.3.1 Architecture
Similar to 2D jellyfish control, we use a 3D U-Net as the backbone of our diffusion model, in both DiffPhyCon-lite and DiffPhyCon methods (detailed in the following subsections). Details of the architecture are same to Table 6.
G.3.2 DiffPhyCon-lite
Except for three fluid features, we use the accumulated proportion of smoke passing through the target exit as another feature. To make this feature (of shape $[T]$), align with fluid features (of shape $[T,3,64,64]$) in shape, we expand it to shape $[T,1,64,64]$ along spatial dimension by value copy. The horizontal and vertical control signals are represented by a tensor of shape $[T,2,64,64]$, where uncontrollable positions are filled with zeros. Then these four state features and control signals are stacked along the channel dimension and we get a tensor of shape $[T,6,64,64]$ as the model input. The model output contains predicted noise of the state features and control signals. Thus its shape is also $[T,6,64,64]$.
G.3.3 DiffPhyCon
The denoising network of $p(\mathbf{w}|\mathbf{c})$ also adopts the 3D U-Net architecture. The input and output size is $[T,2,64,64]$.
G.4 Training, Inference, and Evaluation
Training. The batch size is chosen as 6. The number of iterations and learning rate schedule are same to 2D jellyfish control. The training is performed on two NVIDIA Tesla A6000 GPUs with 48 GB memory for about 2 days.
Inference. The pipeline of inference is similar to 2D jellyfish control as shown in Figure 13. The difference is that we do not involve any surrogate model in this task. In each sampling step, we use the estimated proportion of smoke failing to exit through the target exit as guidance. The inference is performed on one NVIDIA Tesla A6000 GPU with 48 GB memory for about 3 hours for 50 testing samples.
Evaluation. The inference outputs the control force $\mathbf{w}_{[0,T]}$ of $T=64$ steps for 50 testing samples. In simulation, for each testing sample, the initial state and the generated control signals are input to the simulator. The simulator outputs the proportion of smoke failing to exit through the target exit for the final time step. Then the proportions are averaged over 50 test samples.
Appendix H 1D Burgers’ Equation Control Baselines
H.1 PID
@@
As PID itself is a SISO control method, ANN-PID uses a neural network to get multiple sets of PID parameters to do multiple SISO PID controls for MIMO control in the Burgers’ equation. But here, ANN-PID requires the dimensions of inputs and outputs to be the same, so it can only cope with full observation, full control control problems, and partial observation, partial control control problems. Besides, the ANN-PID controller has 2 training setups, including directly interacting with the solver and interacting with the 1D surrogate model in Appendix J.
H.2 SAC
@@
H.3 Supervised Learning
The paper [24] proposes a supervised-learning-based control algorithm that takes a neural operator as a surrogate model to solve control problems. It contains two stages. In the first stage, we take a neural operator to learn the physical constraint as Appendix J. The three CNNs respectively reconstruct $u$, reconstruct $\mathbf{w}$ and learn the transition from $u_{t}$ to $u_{t+1}$. More details are in Appendix J. In the second stage, these three neural networks are used as surrogate models to calculate the gradient of the objective function with respect to the control input. We consider the control $\mathbf{w}$ as a learnable parameter and update it with the gradient.
@@ H.5 BPPO
Appendix I 2D Jellyfish and Smoke Control Baselines
I.1 MPC and SL
Model Predictive Control (MPC) [57] is a control strategy that solves an optimization problem repeatedly to determine the optimal control inputs for a dynamic system. It operates over a finite prediction horizon, optimizing a cost function and applying only the first control action. In the 2D jellyfish movement control problem, MPC uses the control sequences $\mathbf{w}$ and the fluid states as internal state variables. Without the need to train a control agent model, MPC relies on 2D surrogate models mentioned in Appendix J to estimate future states based on the current state and control. We use backpropagation to compute the gradient, update the control action sequences, and optimize the control objective $\mathcal{J}$ in Eq. (19). For every time step, we get the optimized sequences from this time step forward in this way, and only the first control sequence of the optimized control sequences is applied. Compared with MPC, the Supervised learning (SL) method [24] only optimizes the entire control sequences and employs the entire sequences.
@@ I.1 MPC and SL
I.2 SAC
For the 2D case, the algorithm and basic architecture of SAC are the same as the 1D case in Appendix H.2. When designing the reward function for the 2D jellyfish movement control, we find the periodic condition of the opening angle curves is hard to constrain. So to satisfy the periodic condition better, we include the distance between $\mathbf{w}_{t}$ of every time step $t$ and $\mathbf{w}_{0}$. Also, we both consider the squared and absolute error of $(\mathbf{w}_{t}-\mathbf{w}_{0})$ since they respectively constrain the periodic condition when $(\mathbf{w}_{t}-\mathbf{w}_{0})$ is large and small. As a result, the reward function of time step $t$, force $F_{t}$, opening angle $(\mathbf{w}_{t-1},\mathbf{w}_{t})$ and condition angle $\mathbf{w}_{0}$ is defined as
@@ I.2 SAC
where $\lambda_{1}$, $\lambda_{2}$ are weights of different terms. As for the 2D incompressible fluid control, we set the reward function as the percentage of smoke passing through the target bucket.
We take the Adam optimizer [29] to update the weights of networks and the temperature parameter as in the 1D experiment. The hyperparameters are reported in Table 15. In particular, we take the best checkpoint to evaluate the final performance, thus the actual number of training episodes for each setting ranges from about 100 to about 300.
@@
I.3 BC
For the 2D task, the algorithm and basic architecture of BC are the same as the 1D case in Appendix H.4. In Behavior Cloning, the policy network is trained using supervised learning techniques, where the objective is to minimize the difference between the predicted actions and the actions taken by the non-expert in the training data. The loss function is typically the mean squared error between the predicted actions and actions in training data. The training data, consisting of state-action pairs, is collected from non-expert random demonstrations and stored in a dataset. In the inference, we evaluated the same objective function as the SAC. The detailed values of hyperparameters are provided in Table 16.
@@
Hyperparameters for 2D control:
@@
I.4 BPPO
For the 2D task, the algorithm and basic architecture of BPPO are the same as the 1D case in Appendix H.5. In jellyfish movement control, when designing the reward function, we find the periodic condition of the opening angle curves is hard to constrain. So to satisfy the periodic condition better, we include the distance between $\mathbf{w}_{t}$ of every time step $t$ and $\mathbf{w}_{0}$. Also, we both consider the squared and absolute error of $(\mathbf{w}_{t}-\mathbf{w}_{0})$ since they respectively constrain the periodic condition when $(\mathbf{w}_{t}-\mathbf{w}_{0})$ is large and small. As a result, the reward function of time step $t$, force $F_{t}$, opening angle $(\mathbf{w}_{t-1},\mathbf{w}_{t})$ and condition angle $\mathbf{w}_{0}$ is defined as
@@ I.4 BPPO
Hyperparameters for 2D control:
@@
Appendix J Surrogate Models
J.1 1D Surrogate Model
@@
J.3 2D Jellyfish Boundary Mask and Offsets Updater
@@
Appendix K Additional Results of Experiments
Table 19: More results of 1D Burgers’s equation control. Italic font denotes unfair results, bold font denotes the best model among fairly evaluated baselines, and underline denotes the second best model among them.
PO-FC FO-PC PO-PC
$\mathcal{J}$downarrow $\mathcal{J}$downarrow $\mathcal{J}$downarrow
SAC (online) 0.01567 0.034092 0.02768
DiffPhyCon-lite 0.01139 0.00037 0.00494
DiffPhyCon 0.01103 0.00037 0.00494
K.1 DiffPhyCon for 1D Burgers’ Equation
Since online RL methods are effective in a wide domain of control tasks, we tested the performance of the classical RL method, SAC (online), as shown in Table 19. The online SAC belongs to a different setting than the offline training setting considered in our work. Even with access to more information from the environment, online SAC is less competitive to our DiffPhyCon in Burgers’ control task.
@@
It can be observed from Figure 16 that there is a clear correlation between $\mathcal{J}$ and $\mathcal{E}$. This is because, the higher the prediction error is, the less accurate the solution to the inverse control problem would be. The correlation proves that our DiffPhyCon learns the dynamics of the system (i.e., the dependency of $\mathbf{u}$ on $\mathbf{w}$), based on which the control sequence is generated. In partial observation settings, the randomness is larger than in the full observation setting, which demonstrates the challenge of controlling partially observed systems. The reference line denotes $\mathcal{J}=\mathcal{E}$ where the control objective completely depends on the prediction accuracy. In Figure 16 (b), samples cluster tightly around the line, which implies DiffPhyCon can find near-optimal control sequence based on the predicted trajectory in the corresponding level of prediction accuracy. This ability is remarkable in that DiffPhyCon does not rely on explicit gradient descent like those in the adjoint method and SL.
K.2 2D Jellyfish Movement
In this subsection, we record extensive results of 2D jellyfish movement control. To further compare baselines with DiffPhyCon in more settings, we change $\zeta$ in $\mathcal{J}$ and results of $\zeta=500$, $\zeta=1000$ and $\zeta=2000$ are listed in Table 20 and Table 21. The results show that DiffPhyCon outperforms other baselines in diverse settings with different objectives.
@@
K.3 Myopic failure modes of SAC
Table 23: Efficiency comparison on 1D Burgers’ equation. Inference time is tested on a Tesla-V100 GPU with 8 CPUs.
Methods Training time (hours) Inference time (seconds)
DiffPhyCon-lite 1.7 (1 A100-80G GPU, 8 CPUs) 21.13
DiffPhyCon 4.4 (1 A100-80G GPU, 8 CPUs) 58.97
DiffPhyCon-DDIM (8 sampling steps) 1.7 (1 A100-80G GPU, 8 CPUs) 0.53
PID 0.1 (1 A100-80G GPU, 8 CPUs) 3.61
SL 2.6 (1 V100-32G GPU, 12 CPUs) 74.85
SAC 10.5 (1 A6000-48G GPU, 16 CPUs) 0.11
BC 8.8 (1 V100-32G GPU, 12 CPUs) 1.22
BPPO 8.9 (1 V100-32G GPU, 12 CPUs) 0.82
Table 24: Efficiency comparison on 2D jellyfish movement. Inference time is tested on a Tesla-V100 GPU with 8 CPUs.
Methods Training time (hours) Inference time (seconds)
DiffPhyCon 62 (2 A100-80G GPUs, 32 CPUs) 252.2
DiffPhyCon-DDIM (8 sampling steps) 62 (2 A100-80G GPUs, 32 CPUs) 12.6
MPC 52.1 (1 A100-80G GPU, 16 CPUs) 1401.7
SL 52.1 (1 A100-80G GPU, 16 CPUs) 133.5
SAC 9.5 (1 A100-80G, 16 CPUs) 0.2
BC 2.8 (1 A100-80G, 16 CPUs) 0.99
BPPO 3.0 (1 A100-80G GPU, 16 CPUs) 1.08
K.4 Efficiency Comparison
We test the training and inference (control) efficiency of DiffPhyCon and baselines on both 1D Burgers’ equation and 2D jellyfish movement tasks. On the 1D Burgers’ equation task, we test the efficiency of partial observation and full control (PO-FC); on the 2D jellyfish movement task, we test the efficiency of full observation (FO).
The results are presented in Table 23 and Table 24. For training efficiency, since models with different sizes are trained on different machines, we report the training time of models along with the machine information. The inference efficiency is tested on a single Tesla-V100 GPU with 8 CPUs and the average inference times over all test samples are reported. For training efficiency, we have the following observations. On 1D Burgers’ equation, our DiffPhyCon is comparable with SL and more efficient than reinforcement learning (RL) methods like SAC, BC, and BPPO. On 2D jellyfish movement, our DiffPhyCon costs similar training time compared with MPC and SAC, but more time than RL methods. For inference efficiency, these results reveal that although DiffPhyCon is not as efficient as RL methods, it has competitive efficiency compared to SL and MPC. In particular, by using the fast sampling method DDIM [60], DiffPhyCon could be accelerated significantly by reducing the number of sampling steps.
Appendix L Effect of Hyperparameters
L.1 Effect of $\gamma$
We tested different scheduling of the prior reweighting $\gamma$ and found their performance similar. Finally, we use the same cosine schedule as that in the DDPM noise for clarity. Our experiment results also revealed that the performance of DiffPhyCon in 1D Burgers equation control is insensitive to the prior reweighting intensity $\gamma$ as we reported in Table 1. Besides, we conducted a more comprehensive experiment on the prior reweighting intensity $\gamma$ in all three settings (FO-PC; PO-FC; PO-PC) as in Table 25, 26, and 27. The results show that prior reweighting does not significantly impact the performance of DiffPhyCon in 1D Burgers control task, where State MSE denotes the deviation between generated $u$ and the ground truth $u_{\text{gt}}$ given by the numerical simulator based on generated $f$.
$\gamma$ 0.0 0.3 0.5 0.7 1.0
$\mathcal{J}_{\text{actual}}$ 0.00038 0.00037 0.00037 0.00037 0.00037
State MSE 0.00063 0.00043 0.00034 0.00028 0.00025
$\gamma$ 0.0 0.3 0.5 0.7 1.0
$\mathcal{J}_{\text{actual}}$ 0.01159 0.01155 0.01152 0.01148 0.01139
State MSE 0.02643 0.02642 0.02642 0.02641 0.02639
$\gamma$ 0.0 0.3 0.5 0.7 1.0
$\mathcal{J}_{\text{actual}}$ 0.00493 0.00493 0.00493 0.00493 0.00494
State MSE 0.00675 0.00667 0.00665 0.00664 0.00663
@@
In such a setting, varying $\lambda$ changes how the generated samples are shifted towards lower energy regions, but an excessively large $\lambda$ causes the sample to deviate from the physically feasible region. Results in Table 30, 31, and 32 show that the behavior of DiffPhyCon under different $\lambda$ is expected. Note that “physically feasible” samples here are samples with lower $\mathcal{J}_{\text{actual}}$.
@@
2D jellyfish movement control. In our jellyfish movement control task, we use explicit guidance with $\lambda$ involved. Since diffusion models denoise gradually, similar to the hyperparameter $\gamma$, we use a varying sequence of $\lambda=$ $\lambda_{k}$$=\lambda_{0}\times$ $\beta_{k}$, where $\beta_{k}$ is the schedule of variances of noise in DDPM [2] increasing monotonically from $\beta_{1}=0.0003$ to $\beta_{K}=1$. The motivation of this schedule is that the confidence of gradient estimation becomes stronger along with the denoising process from $k=K$ to $k=1$. The results are shown in the Table 33. We can observe that as $\lambda_{0}$ (which determines the $\lambda$ schedule) increases from 0 to 0.6, the performance of DiffPhyCon improves consistently. Compared to Table 3, DiffPhyCon always outperforms baselines for a wide range of $0.1\leq\lambda_{0}\leq 0.5$. When $\lambda_{0}\geq 0.6$, the results are no longer valid since the generated control sequences may be unfeasible. This reflects that $\lambda$ should not be overly large to make it safe, which is a consistent conclusion with our previous analysis.
@@
Appendix M Extensions to Finer-grained Jellyfish Movement Control Task
@@
M.1 Experimental Setting
Similar to previous setups in Appendix F, this extension aims to control the movement of a flapping jellyfish with two wings in a 2D fluid field where fluid flows with constant initial speed. However, in this task, we parametrize the jellyfish’s boundary as the coordinate change ($\Delta x$, $\Delta y$) for each cell within the boundary, which serves as the control sequence $\mathbf{w}$. Undoubtedly, this is a high-dimensional and complex control task. Firstly, the control sequence $\mathbf{w}$ is elevated to three dimensions. In addition to ensuring optimization towards the control objective, maintaining consistency in the movement of different cells within the jellyfish boundary is also crucial. Bearing this in mind, we have chosen to employ DiffConPDE-lite for this experimental setup to minimize potential disruptions to the jellyfish boundary. Regarding the PDE state, the experiment adopts the full observation setup. All other settings remain the same with Appendix F.
M.2 Data Generation
@@
M.4 Training, Inference, and Evaluation
@@
Appendix N Limitation and Future Work
There are still several limitations of DiffPhyCon that inspire future work. Firstly, our approach is data-driven, so it is not guaranteed to achieve optimal solutions and also lacks estimation of the error gap between the generated control sequence and the optimal ones. Secondly, the training of DiffPhyCon is currently conducted in an offline fashion, without interaction with a ground-truth solver. Incorporating solvers into the training framework could facilitate real-time feedback, enabling the model to adapt dynamically to the environment and discover novel strategies and solutions. Furthermore, our proposed DiffPhyCon presently operates in an open-loop manner, as it does not consider real-time feedback from solvers. Integrating such feedback would empower the algorithm to adjust its control decisions based on the evolving state of the environment.
Appendix O Social Impact Statements